OpenStudy (eric_d):

The expression 2^3+ax^2+b can be divided exactly by (x+1), & its remainder is 16 when divided by (x-3). find values of a & b @ganeshie8

ganeshie8 (ganeshie8):

use below : if (x-k) is a factor of a f(x), then f(k) = 0

ganeshie8 (ganeshie8):

since (x+1) is a factor of f(x), f(-1) = 0

OpenStudy (eric_d):

2(-1)^3+a(-1)^2+b=0 like this

ganeshie8 (ganeshie8):

yes, simplify.. that gives you first equation

OpenStudy (eric_d):

-2+a+b=0

ganeshie8 (ganeshie8):

or a+b = 2

OpenStudy (eric_d):

b=2-a

OpenStudy (eric_d):

ok..

ganeshie8 (ganeshie8):

use the other piece of given information to get the second equation : ` its remainder is 16 when divided by (x-3).`

OpenStudy (eric_d):

x=3 2(3)^3+a(3)^2+b=16 b=-38-9a.........2 like this

ganeshie8 (ganeshie8):

yep!

ganeshie8 (ganeshie8):

two equations and two unknowns you can solve them

OpenStudy (eric_d):

Okay Nxt question:

OpenStudy (eric_d):

find value of k if (x+1) is a factor of 2x^3+7x^2+kx-3.Using this value k,solve the equation mentioned sorry 4 d delay

OpenStudy (anonymous):

Again.. as X+1 is a factor..means X+1=0..putting this value in the eqn...find the value of K

OpenStudy (eric_d):

Okay.Understood!

OpenStudy (anonymous):

When this is done.. U will end up with a eqn with the absence of K.. but the presence of its value...Divide this eqn by X+1 as it is a factor....

OpenStudy (eric_d):

Ok..Thnx

OpenStudy (anonymous):

U will have a QUADRATIC EQUATION..U are talented enough t factorize it. Rite ?!! :P :D

OpenStudy (eric_d):

ys

OpenStudy (eric_d):

Nxt question: The expression x^3-4x^2+x+6 & x^3-3x^2+2x+k have common factor.Find possible values of k

OpenStudy (anonymous):

This Is Tricky

OpenStudy (mathmale):

@eric_d: Please present this "Nxt question" as a new question (in the "Ask a question" box, not as something added on to a previous discussion. Thanks.

OpenStudy (eric_d):

Done....!

OpenStudy (eric_d):

I hd re post the question...